# Bigraded path homology and the magnitude-path spectral sequence

### 27 Mar 2026 v2

@@ 6.3. Proof of the excision theorem
Suppose we are in the situation of Definition 6.12, and that $a=\pi(x)$ and $\ell=d(a,x)$. Then the homology of $A_{\ast,\ell}(a,x)$ is a single copy of $R$ in degree $1$, generated by the homology class of the tuple $(\pi(x),x)$.
@@
This quotient has a basis consisting of the tuples $(x_{0},\ldots,x_{k})$ of vertices of $X$ that have finite length and that do not lie entirely in $A$ or entirely in $X\setminus A$. Since $A\hookrightarrow X$ is a cofibration, there are no paths from vertices of $X\setminus A$ into $A$, and so any such tuple satisfies $x_{0},\ldots,x_{i}\in A$ and $x_{i+1},\ldots,x_{k}\in X\setminus A$ for some $i$ in the range $0\leq i<k$.
@@
by setting $F_{i}$ to be the span of all those tuples $(x_{0},\ldots,x_{k})$ for which $x_{0},\ldots,x_{k-i}\in A$. In other words, the final $i$ entries of a tuple in $F_{i}$ are allowed to be outside $A$, but no more.
@@
Since $F_{i}$ and $F_{i-1}$ are spans of generators, with those for the latter being a subset of those for the former, the quotient $F_{i}/F_{i-1}$ has basis consisting of all those generators that lie in $F_{i}$ but not $F_{i-1}$. These are precisely the tuples for which $x_{0},\ldots,x_{k-i}$ lie in $A$ (so that the tuple is in $F_{i}$), while $x_{k-i+1}$ and all subsequent entries of the tuple do not (so that the tuple does not lie in $F_{i-1}$). So we may write an arbitrary generator of $F_{i}/F_{i-1}$ in the form $(x_{0},\ldots,x_{k-i},z_{1},\ldots,z_{i})$ where $x_{0},\ldots,x_{k-i}\in A$ and $z_{1},\ldots,z_{i}\in X\setminus A$.
The boundary map on $F_{i}/F_{i-1}$ is, as usual, given by the alternating sum of all ways of removing an element from a tuple, ignoring any terms for which the length is decreased. If, in a tuple $(x_{0},\ldots,x_{k-i},z_{1},\ldots,z_{i})$, it is one of the $x_{j}$ that is removed, then we find that we still have the first ki1=(k1)ik-i-1=(k-1)-i terms in $A$. On the other hand, if one of the $z_{j}$ is removed, then in this new tuple the first ki=(k1)(i1)k-i=(k-1)-(i-1) terms lie in $A$, so that the tuple lies in $F_{i-1}$ and therefore vanishes in $F_{i}/F_{i-1}$. From this, we see that the differential of $F_{i}/F_{i-1}$ is given by the alternating sum of all ways to delete one of the $x_{j}$ from a tuple $(x_{0},\ldots,x_{k-i},z_{1},\ldots,z_{i})$, omitting any terms where the length is decreased.
@@ 8. Directed cycles
Observe that each $Z_{m}$ has diameter m1m-1, so that all $Z_{m}$ are long homotopy equivalent to one another and to the singleton, and in particular they all have trivial reachability homology. In contrast, when we consider $1$-homotopy, it is not difficult to see that for $m\geq 2$ no two of the $Z_{m}$ are $1$-homotopy equivalent (when $m<n$ the only maps $Z_{m}\to Z_{n}$ are constant, and for $n>2$ the constant maps $Z_{n}\to Z_{n}$ are only $1$-homotopic to other constant maps). Nevertheless, for $m\geq 3$ the $Z_{m}$ all have the same path homology, namely a single copy of $R$ in degrees 0 and $1$. So path homology cannot distinguish the different oriented cycles. In contrast, we will see that bigraded path homology can distinguish all except for $Z_{1}$ and $Z_{2}$.
@@
Let $m\geq 3$. Then the magnitude homology $\mathrm{MH}_{\ast,\ast}(Z_{m})$ and bigraded path homology $\mathrm{PH}_{\ast,\ast}(Z_{m})$ are both concentrated in bidegrees of the form $(2i,mi)$ and $(2i+1,mi+1)$, in each of which they are free of rank $m$ and rank $1$ respectively. Moreover, the MPSS of $Z_{m}$ satisfies $E^{2}(Z_{m})=\cdots=E^{m-1}(Z_{m})$ while $E^{m}(Z_{m})$ is trivial, consisting of a single copy of $R$ in bidegree $(0,0)$.
@@ 8.1. The complex of ordered partitions with upper bound
Given $\ell,m\in\mathbb{Z}$ with $m\geq 2$, we let $\mathcal{OP}_{\ast}(\ell,m)$ be the chain complex of $R$-modules defined as follows. In degree $k$, $\mathcal{OP}_{k}(\ell,m)$ has basis given by the ordered partitions $(a_{1},\ldots,a_{k})$ of $\ell$ such that $a_{i}<m$ for all $i$. The differential $d$ vanishes for $k\leq 1$, and is given in degrees $k\geq 2$ by summing adjacent entries,
@@
We will see that, modulo a degree shift of $2$, $H_{\ast}(\mathcal{OP}(\ell,m))$ is in fact periodic in $\ell$ with period $m$. This will allow us to prove the theorem by induction. We begin with the base cases.
@@
  • For $\ell=1$, $H_{\ast}(\mathcal{OP}(\ell,m))$ is a single copy of $R$ concentrated in degree $1$, generated by the class of the word $(1)$.
@@ 8.3. Bigraded path homology
In particular, $\mathrm{PH}_{\ast,\ast}(Z_{m})$ is free of rank $1$ in bidegrees of form $(2i,mi)$ and $(2i+1,mi+1)$ for $i\geq 0$, and vanishes in all other bidegrees.
@@ 9. Bi-directed cycles
Let $m,n\geq 1$ and assume further that $\mathfrak{m}=\max(m,n)\geq 3$. Then the bigraded path homology $\mathrm{PH}_{\ast,\ast}(C_{m,n})$ is concentrated in bidegrees $(0,0)$, $(1,0)$ and $(\mathfrak{m},-(\mathfrak{m}-2))$, in each of which it is free of rank $1$. Moreover, the MPSS of $C_{m,n}$ satisfies $E^{2}(C_{m,n})=\cdots=E^{\mathfrak{m}-1}(C_{m,n})$ while $E^{\mathfrak{m}}(C_{m,n})$ is trivial, consisting of a single copy of $R$ in bidegree $(0,0)$. If $\mathfrak{m}=1,2$ then the bigraded path homology is concentrated in bidegree $(0,0)$, where it is free of rank $1$.
@@
whose upper map collapses $A_{m,n}$ to $A_{m,1}$, the latter being a single vertex. The maps $\mathrm{PH}_{\ast,\ast}(A_{m,n})\to\mathrm{PH}_{\ast,\ast}(A_{m,1})$ are isomorphisms since $A_{m,n}\to A_{m,1}$ is a $1$-homotopy equivalence. And the maps $\mathrm{PH}_{\ast,\ast}(C_{m,n},A_{m,n})\to\mathrm{PH}_{\ast,\ast}(C_{m,1},A_{m,1})$ are isomorphisms by the excision theorem, which applies since the vertical maps of our pushout are cofibrations. Then the long exact sequences of the pairs $(C_{m,n},A_{m,n})$ and $(C_{m,1},A_{m,1})$ are related by a commutative ladder, and the last sentence allows us to apply the five lemma to this ladder to obtain the result. ∎

### 27 Apr 2026 v3

@@ 6.3. Proof of the excision theorem
Suppose we are in the situation of Definition 6.12, and that $a=\pi(x)$ and $\ell=d(a,x)$. Then the homology of $A_{\ast,\ell}(a,x)$ is a single copy of $R$ in degree 11, generated by the homology class of the tuple $(\pi(x),x)$.
@@
This quotient has a basis consisting of the tuples $(x_{0},\ldots,x_{k})$ of vertices of $X$ that have finite length and that do not lie entirely in AA or entirely in $X\setminus A$. Since $A\hookrightarrow X$ is a cofibration, there are no paths from vertices of $X\setminus A$ into AA, and so any such tuple satisfies $x_{0},\ldots,x_{i}\in A$ and $x_{i+1},\ldots,x_{k}\in X\setminus A$ for some $i$ in the range $0\leq i<k$.
@@
by setting $F_{i}$ to be the span of all those tuples $(x_{0},\ldots,x_{k})$ for which $x_{0},\ldots,x_{k-i}\in A$. In other words, the final $i$ entries of a tuple in $F_{i}$ are allowed to be outside AA, but no more.
@@
Since $F_{i}$ and $F_{i-1}$ are spans of generators, with those for the latter being a subset of those for the former, the quotient $F_{i}/F_{i-1}$ has basis consisting of all those generators that lie in $F_{i}$ but not $F_{i-1}$. These are precisely the tuples for which $x_{0},\ldots,x_{k-i}$ lie in AA (so that the tuple is in $F_{i}$), while $x_{k-i+1}$ and all subsequent entries of the tuple do not (so that the tuple does not lie in $F_{i-1}$). So we may write an arbitrary generator of $F_{i}/F_{i-1}$ in the form $(x_{0},\ldots,x_{k-i},z_{1},\ldots,z_{i})$ where $x_{0},\ldots,x_{k-i}\in A$ and $z_{1},\ldots,z_{i}\in X\setminus A$.
The boundary map on $F_{i}/F_{i-1}$ is, as usual, given by the alternating sum of all ways of removing an element from a tuple, ignoring any terms for which the length is decreased. If, in a tuple $(x_{0},\ldots,x_{k-i},z_{1},\ldots,z_{i})$, it is one of the $x_{j}$ that is removed, then we find that we still have the first ki1=(k1)ik-i-1=(k-1)-i terms in AA. On the other hand, if one of the $z_{j}$ is removed, then in this new tuple the first ki=(k1)(i1)k-i=(k-1)-(i-1) terms lie in AA, so that the tuple lies in $F_{i-1}$ and therefore vanishes in $F_{i}/F_{i-1}$. From this, we see that the differential of $F_{i}/F_{i-1}$ is given by the alternating sum of all ways to delete one of the $x_{j}$ from a tuple $(x_{0},\ldots,x_{k-i},z_{1},\ldots,z_{i})$, omitting any terms where the length is decreased.
@@ 8. Directed cycles
Observe that each $Z_{m}$ has diameter m1m-1, so that all $Z_{m}$ are long homotopy equivalent to one another and to the singleton, and in particular they all have trivial reachability homology. In contrast, when we consider 11-homotopy, it is not difficult to see that for $m\geq 2$ no two of the $Z_{m}$ are 11-homotopy equivalent (when $m<n$ the only maps $Z_{m}\to Z_{n}$ are constant, and for $n>2$ the constant maps $Z_{n}\to Z_{n}$ are only 11-homotopic to other constant maps). Nevertheless, for $m\geq 3$ the $Z_{m}$ all have the same path homology, namely a single copy of $R$ in degrees 0 and 11. So path homology cannot distinguish the different oriented cycles. In contrast, we will see that bigraded path homology can distinguish all except for $Z_{1}$ and $Z_{2}$.
@@
Let $m\geq 3$. Then the magnitude homology $\mathrm{MH}_{\ast,\ast}(Z_{m})$ and bigraded path homology $\mathrm{PH}_{\ast,\ast}(Z_{m})$ are both concentrated in bidegrees of the form $(2i,mi)$ and $(2i+1,mi+1)$, in each of which they are free of rank $m$ and rank 11 respectively. Moreover, the MPSS of $Z_{m}$ satisfies $E^{2}(Z_{m})=\cdots=E^{m-1}(Z_{m})$ while $E^{m}(Z_{m})$ is trivial, consisting of a single copy of $R$ in bidegree $(0,0)$.
@@ 8.1. The complex of ordered partitions with upper bound
Given $\ell,m\in\mathbb{Z}$ with $m\geq 2$, we let $\mathcal{OP}_{\ast}(\ell,m)$ be the chain complex of $R$-modules defined as follows. In degree $k$, $\mathcal{OP}_{k}(\ell,m)$ has basis given by the ordered partitions $(a_{1},\ldots,a_{k})$ of $\ell$ such that $a_{i}<m$ for all ii. The differential $d$ vanishes for $k\leq 1$, and is given in degrees $k\geq 2$ by summing adjacent entries,
@@
We will see that, modulo a degree shift of 22, $H_{\ast}(\mathcal{OP}(\ell,m))$ is in fact periodic in $\ell$ with period $m$. This will allow us to prove the theorem by induction. We begin with the base cases.
@@
  • For $\ell=1$, $H_{\ast}(\mathcal{OP}(\ell,m))$ is a single copy of $R$ concentrated in degree 11, generated by the class of the word $(1)$.
@@ 8.3. Bigraded path homology
In particular, $\mathrm{PH}_{\ast,\ast}(Z_{m})$ is free of rank 11 in bidegrees of form $(2i,mi)$ and $(2i+1,mi+1)$ for $i\geq 0$, and vanishes in all other bidegrees.
@@ 9. Bi-directed cycles
Let $m,n\geq 1$ and assume further that $\mathfrak{m}=\max(m,n)\geq 3$. Then the bigraded path homology $\mathrm{PH}_{\ast,\ast}(C_{m,n})$ is concentrated in bidegrees $(0,0)$, $(1,0)$ and $(\mathfrak{m},-(\mathfrak{m}-2))$, in each of which it is free of rank 11. Moreover, the MPSS of $C_{m,n}$ satisfies $E^{2}(C_{m,n})=\cdots=E^{\mathfrak{m}-1}(C_{m,n})$ while $E^{\mathfrak{m}}(C_{m,n})$ is trivial, consisting of a single copy of $R$ in bidegree $(0,0)$. If $\mathfrak{m}=1,2$ then the bigraded path homology is concentrated in bidegree $(0,0)$, where it is free of rank 11.
@@
whose upper map collapses $A_{m,n}$ to $A_{m,1}$, the latter being a single vertex. The maps $\mathrm{PH}_{\ast,\ast}(A_{m,n})\to\mathrm{PH}_{\ast,\ast}(A_{m,1})$ are isomorphisms since $A_{m,n}\to A_{m,1}$ is a 11-homotopy equivalence. And the maps $\mathrm{PH}_{\ast,\ast}(C_{m,n},A_{m,n})\to\mathrm{PH}_{\ast,\ast}(C_{m,1},A_{m,1})$ are isomorphisms by the excision theorem, which applies since the vertical maps of our pushout are cofibrations. Then the long exact sequences of the pairs $(C_{m,n},A_{m,n})$ and $(C_{m,1},A_{m,1})$ are related by a commutative ladder, and the last sentence allows us to apply the five lemma to this ladder to obtain the result. ∎